OpenStudy (anonymous):

You want to have $100,000 when you retire in 30 years. If you can earn 4% interest compounded monthly, how much would you need to deposit now into the account to reach your retirement goal?

OpenStudy (anonymous):

ok, I can help you

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

first, set up the equation

OpenStudy (anonymous):

100000=a(0.04)^30

OpenStudy (anonymous):

y=ab^x

OpenStudy (anonymous):

scratch that, sorry I'm tired

OpenStudy (anonymous):

its alright haha

OpenStudy (anonymous):

yeah you are right, 10,000=ab^x

OpenStudy (anonymous):

ok we want a growth of 0.04 or 4%

OpenStudy (anonymous):

whats next

OpenStudy (anonymous):

1.04

OpenStudy (anonymous):

This will substitute b

OpenStudy (anonymous):

100000=a(1.04)^30

OpenStudy (anonymous):

NICE. You read my mind

OpenStudy (anonymous):

Now what

jimthompson5910 (jim_thompson5910):

that equation would work if you are compounding annually

OpenStudy (anonymous):

now just work backwards (I think (I don't know it might be the sleep))

OpenStudy (anonymous):

we dont have a yet so we dont do that yet

OpenStudy (anonymous):

how do I find A??

OpenStudy (anonymous):

actually @jim_thompson5910 is right

OpenStudy (anonymous):

it's a monthly payment, and you are using the year

OpenStudy (anonymous):

so we need to find how many months are in 30 years

OpenStudy (anonymous):

100000=a(1.48)^360

OpenStudy (anonymous):

there now how do i find a

OpenStudy (anonymous):

nice now we need to divide 10,000 by 1.04^360 to find a

jimthompson5910 (jim_thompson5910):

4% annual rate 4/12 = 0.333% monthly rate so instead of 1.04 it should be 1.003333

OpenStudy (anonymous):

... wow, I'm very tired. I can't believe I missed that. He's right *Slaps self*

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

wait @jim_thompson5910 wouldn't it be 33.3% not 0.333%

OpenStudy (anonymous):

which i dont think you're right jim, you are suppose to do .04/12

OpenStudy (anonymous):

then It would be 3.33%

jimthompson5910 (jim_thompson5910):

(4%)/12 = 0.33333% the 4% will get smaller and it won't increase to 33.33%

jimthompson5910 (jim_thompson5910):

or yeah you could do 0.04/12 = 0.00333...

OpenStudy (anonymous):

ok nvm. So you have 10,000 / 1.00333^360 = A

OpenStudy (anonymous):

100000=a(1.003333)^360

OpenStudy (anonymous):

how do i find a

jimthompson5910 (jim_thompson5910):

divide both sides by (1.003333)^360

OpenStudy (anonymous):

100000/3.3131?

OpenStudy (anonymous):

that's what I said, "10,000 / 1.00333^360"

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah you just compute 100000/3.3131 now

OpenStudy (anonymous):

I got 30183.2121 but thats wrong

jimthompson5910 (jim_thompson5910):

did you round to the nearest penny?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its actually 30179.59, you did something wrong i dunno

jimthompson5910 (jim_thompson5910):

ok try typing in (1+0.04/12)^(360) as the denominator to ensure there is no rounding error

jimthompson5910 (jim_thompson5910):

use a calculator like this one http://web2.0calc.com/ if you don't have a calculator like a TI83

jimthompson5910 (jim_thompson5910):

http://web2.0calc.com/#100000/(1+0.04/12)^(360)

jimthompson5910 (jim_thompson5910):

when you click the link, hit the "=" button on the calculator to have it compute the answer

OpenStudy (anonymous):

Hey guys I'm back, I had to go to the bathroom, and then help my dad, sorry.

OpenStudy (anonymous):

I'm constantly getting 3021.57 on the Bing calculator. I'm going to double check

OpenStudy (anonymous):

When I insert it into the equation, I get 9999.99889

OpenStudy (anonymous):

I think it would be safe to round it, so my answer would be 3021.57 for your question.